OpenStudy (anonymous):

Help:( Have to submit homework in 25 minutes! In a triangle ABC, a = 62 cm, b = 62 cm, and A = 71°. The length of c is?

OpenStudy (phi):

looks like you have an isosceles triangle, so angle B= angle A now you can find angle C now you know enough to use the Law of sines

OpenStudy (campbell_st):

draw a diagram  the triangle is isosceles and angle C = 38 by angle sum of a triangle use the cosine rule \[c^2 = 62^2 + 62^2 - 2\times 62 \times 62 \times \cos(38)\]